Solve for x: (e^x)+(e^x)= 0
Found 2 solutions by jim_thompson5910, Alan3354:
Answer by jim_thompson5910(35256)

(e^x)+(e^x)= 0

2e^x= 0

e^x= 0/2

e^x= 0

ln(e^x) = ln(0)

Since you cannot take the natural log of 0, there are no solutions.

Answer by Alan3354(69443)

Solve for x: (e^x)+(e^x)= 0
-------------
2e^x = 0
e^x = 0
No solution.